# Realigning Incentives to Build Better Software: a Holistic Approach to Vendor Accountability

### v1

@@ Abstract
In this paper, we ask the question of why the quality of commercial software, in terms of security and safety, does not measure up to that of other (durable) consumer goods we have come to expect. We examine this question through the lens of incentives. We argue that the challenge around better quality software is due in no small part to a sequence of misaligned incentives, the most critical of which being that the harm caused by software problems is by and large shouldered by consumers, not developers. This lack of liability means software vendors have every incentive to rush low-quality software onto the market and no incentive to enhance quality control. Within this context, this paper outlines a holistic technical and policy framework we believe is needed to incentivize better and more secure software development. At the heart of the incentive realignment is the concept of software liability. This framework touches on various components, including legal, technical, and financial, that are needed for software liability to work in practice; some currently exist, some will need to be re-imagined or established. This is primarily a market-driven approach that emphasizes voluntary participation but highlights the role appropriate regulation can play. We connect and contrast this with the EU legal environment and discuss what this framework means for open-source software (OSS) development and emerging AI risks. Moreover, we present a CrowdStrike case study complete with a what-if analysis had our proposed framework been in effect. Our intention is very much to stimulate a robust conversation among both researchers and practitioners.
@@ 1.1 Background and Motivation
We contend that there are fundamental problems with the software development and deployment ecosystem and that these problems will only grow with time unless fundamental changes are made. Globally, we interact with more software applications across more consumer and commercial devices that operate and influence more critical parts of our lives. Over time, these software products have been expanding in size and becoming more complex, in part from referencing and incorporating more 3rd party libraries and dependencies. Every additional library makes it more and more difficult to identify design, logic, and implementation vulnerabilities. Indeed, firms may well have no idea about existing or newly discovered vulnerabilities in the third-party code they use and deploy. For example, the widely discussed 2021 log4j incident left companies scrambling to assess whether their own systems are vulnerable owing to the popularity and large-scale reuse of the log4j package via the software supply chain [1]. The ineffectiveness of the lengthy assessment enabled attackers to continue exploiting the Log4Shell vulnerability throughout the next year. With nearly 30,0003000030,00030 , 000 vulnerabilities disclosed in 2024, and increasing monotonically year over year, such incidents are bound to happen even more frequently [2].
@@
Many scholars have argued that the current market does not enforce sufficient accountability on the software or hardware vendor [4, 5, 6, 7]. Indeed, many residential and commercial devices are sold without appreciation for security and privacy concerns or often without any expectation of providing fixes for their software as new vulnerabilities are disclosed. At the same time, the current U.S. legal regime is not capable of holding vendors liable for any losses from faulty or improperly maintained software, economic or otherwise. Consequently, harms are passed downward to end-users and consumers.
Different efforts have attempted to mitigate challenges with faulty software. For example, vulnerability disclosure and bug bounty programs have emerged as effective tools for crowdsourcing vulnerability hunting to security researchers, enabling subsequent patch development by software vendors and, finally, patching by system administrators. As successful as these efforts have been, they only address the symptoms rather than provide cures to the problem. In addition, NIST, BSA, Microsoft, and others have developed secure coding standards (e.g., [8, 9, 10, 11]), which could presumably prevent many of the problems, to begin with; however, they are not enforced and therefore are not systematically employed. Worse, firms often deliberately eschew such best practices in order to cut costs and minimize time to market.
Collectively, we argue that this expanding software universe, combined with difficulties in managing the growing software codebase, the lack of accountability and liability, and partial solutions that address symptoms rather than root causes, together create more opportunities for both accidental failures and malicious attacks, increasing the risk caused by these software components.
@@ 1.2 Misaligned incentives and proposed realignment
And so, how can we produce “better” software? In this paper we examine this question through the lens of incentives. We argue that the challenge around better software is due in no small part to a sequence of misaligned incentives. Reliability, safety, and security features in a software product can appear invisible to the user and, thus, hard for the software developer to monetize. Therefore, spending resources on improving these features may never lead to immediate or substantial returns. Moreover, the first-mover advantage incentivizes developers and vendors to release (new versions of) products quickly without fixing (sometimes known) vulnerabilities.
@@
While many of the elements in the proposed ecosystem currently exist, others may need to be established or re-imagined. Some address ex-ante issues (what happens prior to the software development or adoption), and some ex-post (what happens when liability issues emerge following the revelation of software vulnerabilities to which business loss can be attributed). Above all, we argue that the incentives among these components need to be (re-)aligned for the ecosystem to function properly. Each component can provide value on its own, but we argue that collectively, they can help create an effective system with the right incentives and checks and balances to induce more vendor responsibility in producing secure software. We discuss each component and the incentive relationships in detail in Section 3.
@@
Importantly, we do not assume that the better alignment of incentives will eradicate all software flaws and software-based cyber incidents. Nor do we advocate for heavy-handed government regulation to oversee and interfere with commercial software development and deployment. Instead we seek to introduce mechanisms in order to bolster accountability, reduce externalities, and balance information asymmetries within the software development and deployment ecosystem.
@@ 2 Current State of Software Regulation
Below we review the developments in the regulatory landscape most relevant to the present paper. A more comprehensive review is provided in Appendix A for the interested reader.
@@
The EU Council is arguably at the forefront of the “product” movement: it enacted a new Product Liability Directive (PLD) [15] that came into force in December 2024, with the intent to explicitly designate software (including AI software) a “product”; this in turn comes with strict liability clauses. The PLD allows individuals, including consumers, to seek compensation from manufacturers on a strict liability basis for defective products and, in some cases, their components in the EU market. Its goal is to simplify the claims process for damages caused by product defects. It expands liability to nearly all supply chain operators and ensures consumer protection regardless of a product’s origin (EU or non-EU). Consumers can seek compensation in complex cases, including safety/security regulation breaches (the draft AI Liability Directive (AILD [16]) further specifies such motions in the context of AI software). The removal of arbitrary thresholds allows full compensation for damages. With a very broad definition of product, this PLD addresses new technologies such as cyber vulnerabilities, essential digital services, and software updates. Software components include any item (e.g., a software library), raw material, or service (e.g., when software initiates remote calls to a SaaS instance) integrated into or connected with a product, though standalone services are generally not covered. Notably, free and open-source software (OSS) outside commercial activity is excluded.
Software as (mostly) a service
@@
In addition to self-attestation, a cybersecurity “safe harbor”, based on the certification of organizations (with a similar approach already codified in the EU NIS2 Directive [20]), has received considerable attention [21]. Proponents argue that a certification safe harbor can (over time) create incentives to improve collective security practices, including supply chain security, via involving certified upstream technology providers. Dempsey further suggests a hybrid approach, combining a product-based “floor” for liability and a process-based “ceiling” for safe harbor [4]. His proposed mechanism for liability is three-fold. First, a rule-based approach would define the minimum standard of care for software focusing on product features and behaviors to be included or avoided (e.g., any weaknesses tracked in the MITRE Common Weakness Enumeration (CWE) database [^2] https://cwe.mitre.org/). Second, as software can be complex and dynamic, a liability regime and analysis are needed to assess non-trivial design and implementation flaws; essentially, a software audit. This approach shares similarities with European legislation that has already codified voluntary [22] and mandatory [23] software audit and certification. Third, to provide a safe harbor to developers who followed secure coding recommendations, the proposal suggests a liability waiver or limitation, which makes the outcome predictable for vendors. Clearly defining the conditions under which a vendor qualifies for safe harbor provides legal certainty, reducing the need for costly and prolonged expert debates over what qualifies as reasonable (non-negligent) conduct [7].
@@ 3 Designing A Better Software Ecosystem
Our proposal first begins in the upper path, establishing a minimum set of software development requirements (the floor) and safe harbor practices (the ceiling). As these policies are adopted, they increase the liability of commercial software developers for their software products. In addition, a set of process and product audits, combined with other measures, increase transparency by signaling software quality to stakeholders. Together, these forces – liability and transparency – combine to increase the accountability of vendors. This, in turn, increases the cost of releasing low-quality (i.e., vulnerable, untested, unmaintained) software, which in turn reduces the vendor’s incentive to rush to release software.
@@ 3.1 Audit and Software Clearinghouse
The audit service should, at a minimum, include the following components. (1) Completing an audit should, in general, be optional/voluntary but may be mandatory in some specific and limited scope; e.g., certain types of software, such as those used in critical infrastructure, may be required to go through an audit. (2) Passing an audit waives the vendor of future liability up to some upper limit for fines or penalties it may face if/when vulnerabilities are discovered or if harm is attributed to its software. (3) The audit process itself should go through regular updates and improvements. Thus, the audit also carries a version number that is included in the certification it issues. (4) Software that fails an audit should not be allowed to be released.
@@
One of the key outcomes of the audit system will be a Software Clearinghouse, a repository of software that has been certified by the auditor, including libraries that feed into a myriad of other software products and systems. In principle, software in the clearinghouse has been certified, and its subsequent use (e.g., as dependencies, clearly articulated in the software’s Bill of Materials (SBOM) [24]) carries limited liability waiver as described in the previous section. We believe this will encourage customers to maximally use certified software and vendors to build on already certified dependencies. Such an approach has the potential to (partially) prevent the challenges and well-documented struggles of customers trying to identify vulnerable dependencies within their own systems upon the disclosure of newly identified vulnerabilities (especially affecting popular libraries/packages), just as it happened in the Log4Shell case [1].
@@
Software product and process audits have their own set of challenges; some even question whether cybersecurity-related audits could be effective [25]. In reality, software quality assurance in general and audit, in particular, is its own field of study [26] with its own standards and recommendations, e.g., IEEE 1028-2008 [27]. In relation to product-based security testing, defining a set of minimum requirements (i.e., a “floor”) can enhance both the effectiveness of the audit and the interpretation of the resulting certification (see Section 3.2 for details).
@@
As for the technical security testing methodology, the widespread DevOps philosophy and the resulting Continuous Integration/Continous Development (CI/CD) approach, designed to streamline and speed up software production and feature delivery pose unique challenges. One such challenge is the need for rapid and continuous security testing, manifesting in the DevSecOps trend. Such speed-up requires automation, which is far from trivial for more powerful testing methods, e.g., dynamic testing [28]. In the context of the proposed Software Clearinghouse, one critical enabling technology will be quick and automated (regression) testing, to keep up with the rapid releases of new versions of already certified software packages without re-doing a full audit [29].
@@ 3.2 Software Liability
This approach can directly lead to a meaningful interpretation of liability, as it delineates the responsibility: the vendor is (largely) not responsible for the residual risk its product carries as long as it adheres to a specified standard of process; it is, however, wholly responsible for meeting a set of minimum quality requirements even if it has adhered to the process standards. From the user’s point of view, this delineation helps them understand where their own responsibilities and risks lie: if a software has passed both the product audit and the process audit, then the user is aware that the vendor can no longer be faulted for harm that may be caused by the software. If, on the other hand, the software passed the product audit but failed the process audit, then the vendor may be responsible for compensating the user for a large share of the harm attributable to the software. In either case, the user is better informed and can use that knowledge to seek the right type of risk protection (see the next section on insurance). In the same vein, this approach can help resolve legal disputes by establishing who is responsible when harm occurs.
@@
As to how to define the minimum quality requirements, i.e., the floor, one may turn to “unforgivable vulnerabilities” [30]. Such weaknesses are trivial to find and occur time and time again, demonstrating a “systematic disregard” for secure development practices. In a recent proposal, the UK National Cyber Security Centre (NCSC) proposed a method to classify vulnerabilities into “forgivable” and “unforgivable”  [31]. The method focuses on finding the root cause of vulnerabilities and their respective top-level mitigations based on the MITRE CWE catalog. The mitigation is then assigned an “ease of implementation” score (taking cost, familiarity, and technical feasibility into consideration); vulnerabilities with “easy” mitigations are then declared unforgivable. Simply put, such weaknesses should not appear in reasonably secure software.
@@
In turn, the process-based ceiling may adopt ideas from the secure software development lifecycle, e.g., Microsoft’s SDL [32]. SDL promotes a DevSecOps approach integrating security into DevOps (DevSecOps). It applies to all software development models and supports diverse software types (firmware, AI, IoT, web services, etc.) across multiple platforms (cloud, on-premises, mobile, SaaS, etc.). The SDL emphasizes 10 key security practices to enhance software security throughout the whole lifecycle. Adding to this, the 2024 White House report on secure and measurable software [33] further promotes the use of advanced secure technologies such as memory-safe programming languages (e.g., Rust, Go, etc.), memory-safe hardware, and formal methods to prove the correctness of the developed code. Furthermore, the report advocates for the use of cybersecurity quality metrics. Fascinatingly, it also acknowledges the need for shifting market forces, i.e., re-aligning incentives, to improve these metrics, which is exactly our objective in this paper.
@@
Interestingly, the floor-ceiling approach can also help settle the two key pending legal issues concerning software liability mentioned earlier: (1) whether software will remain a “service” in the eyes of the law or will start being viewed as a “product” at some point; and (2) whether legal protections against lawsuits should be based on satisfactory process, product quality, or both. As far as we know, there have been no direct liability lawsuits until very recently (see Delta Airlines vs. CrowdStrike in our case study in Section 4. This is directly tied to the fact that there is a lack of more precise definitions of liability, which is much needed. As a reference, the EU GDPR was able to define both duty and standard of care, leading to substantial fines for mishandling personal data and prompting a reassessment of cybersecurity investments (as risk-matching cybersecurity countermeasures have also been prescribed for safeguarding personal data). This gives one hope that similar definitions may emerge for the software industry.

### v2

@@ Abstract
In this paper, we ask the question of why the quality of commercial software, in terms of security and safety, does not measure up to that of other (durable) consumer goods we have come to expect. We examine this question through the lens of incentives. We argue that the challenge around better quality software is due in no small part to a sequence of misaligned incentives, the most critical of which being that the harm caused by software problems is, by and large, shouldered by consumers, not developers. This lack of liability means software vendors have every incentive to rush low-quality software onto the market and no incentive to enhance quality control. Within this context, this paper outlines a holistic technical and policy framework we believe is needed to incentivize better and more secure software development. At the heart of the incentive realignment is the concept of software liability. This framework touches on various components, including legal, technical, and financial, that are needed for software liability to work in practice; some currently exist, some will need to be re-imagined or established. This is primarily a market-driven approach that emphasizes voluntary participation but highlights the role appropriate regulation can play. We connect and contrast this with the EU legal environment and discuss what this framework means for open-source software (OSS) development and emerging AI risks. Moreover, we present a CrowdStrike case study complete with a what-if analysis had our proposed framework been in effect. Our intention is very much to stimulate a robust conversation among both researchers and practitioners.
@@ 1.1 Background and Motivation
We contend that there are fundamental problems with the software development and deployment ecosystem and that these problems will only grow with time unless fundamental changes are made. Globally, we interact with more software applications across more consumer and commercial devices that operate and influence more critical parts of our lives. Over time, these software products have been expanding in size and becoming more complex, in part from referencing and incorporating more third-party libraries and dependencies. Every additional library makes it more and more difficult to identify design, logic, and implementation vulnerabilities. Indeed, firms may well have no idea about existing or newly discovered vulnerabilities in the third-party code they use and deploy. For example, the widely discussed 2021 log4j incident left companies scrambling to assess whether their own systems are vulnerable owing to the popularity and large-scale reuse of the log4j package via the software supply chain [1]. The ineffectiveness of the lengthy assessment enabled attackers to continue exploiting the Log4Shell vulnerability throughout the next year. With nearly 30,0003000030,00030 , 000 vulnerabilities disclosed in 2024, and increasing monotonically year over year, such incidents are bound to happen even more frequently [2].
@@
Many scholars have argued that the current market does not enforce sufficient accountability on the software or hardware vendor [4, 5, 6, 7]. Indeed, many residential and commercial devices are sold without appreciation for security and privacy concerns, or often without any expectation of providing fixes for their software as new vulnerabilities are disclosed. At the same time, the current U.S. legal regime is not capable of holding vendors liable for any losses from faulty or improperly maintained software, economic or otherwise. Consequently, harms are passed downward to end-users and consumers.
Different efforts have attempted to mitigate challenges with faulty software. For example, vulnerability disclosure [8] and bug bounty programs [9] have emerged as effective tools for crowdsourcing vulnerability hunting to security researchers, enabling subsequent patch development by software vendors and, finally, patching by system administrators. As successful as these efforts have been, they only address the symptoms rather than provide a cure to the problem. In addition, NIST, BSA, Microsoft, and others have developed secure coding standards (e.g., [10, 11, 12, 13]), which could presumably prevent many of the problems, to begin with; however, they are not enforced and therefore are not systematically employed. Worse, firms often deliberately eschew such best practices in order to cut costs and minimize time to market.
Collectively, we argue that this expanding software universe, combined with difficulties in managing the growing software codebase, the lack of accountability and liability, and partial solutions that address symptoms rather than root causes, together create more opportunities for both accidental failures and malicious attacks, increasing the risk caused by these software components. The importance and non-triviality of this issue and its policy context have been recognized by the broader research community, as illustrated by its inclusion in the top 10 “cyber hard problems” compiled by the U.S. National Academy of Sciences [14] (Problem 5).
@@ 1.2 Misaligned incentives and proposed realignment
And so, how can we produce “better” software? In this paper, we examine this question through the lens of incentives. We argue that the challenge around better software is due in no small part to a sequence of misaligned incentives. Reliability, safety, and security features in a software product can appear invisible to the user and, thus, hard for the software developer to monetize. Therefore, spending resources on improving these features may never lead to immediate or substantial returns. Moreover, the first-mover advantage incentivizes developers and vendors to release (new versions of) products quickly without fixing (sometimes known) vulnerabilities.
@@
While many of the elements in the proposed ecosystem currently exist, others may need to be established or re-imagined. Some address ex-ante issues (what happens prior to the software development or adoption), and some ex-post (what happens when liability issues emerge following the revelation of software vulnerabilities to which business loss can be attributed). Above all, we argue that the incentives among these components need to be (re)aligned for the ecosystem to function properly. Each component can provide value on its own, but we argue that collectively, they can help create an effective system with the right incentives and checks and balances to induce more vendor responsibility in producing secure software. We discuss each component and the incentive relationships in detail in Section 3.
@@
Importantly, we do not assume that the better alignment of incentives will eradicate all software flaws and software-based cyber incidents. Nor do we advocate for heavy-handed government regulation to oversee and interfere with commercial software development and deployment. Instead, we seek to introduce mechanisms in order to bolster accountability, reduce externalities, and balance information asymmetries within the software development and deployment ecosystem.
@@ 2 Current State of Software Regulation
Below, we review the developments in the regulatory landscape most relevant to the present paper. A more comprehensive review is provided in Appendix A for the interested reader.
@@
The EU Council is arguably at the forefront of the “product” movement: it enacted a new Product Liability Directive (PLD) [18] that came into force in December 2024, with the intent to explicitly designate software (including AI software) a “product”; this in turn comes with strict liability clauses. The PLD allows individuals, including consumers, to seek compensation from manufacturers on a strict liability basis for defective products and, in some cases, their components in the EU market. Its goal is to simplify the claims process for damages caused by product defects. It expands liability to nearly all supply chain operators and ensures consumer protection regardless of a product’s origin (EU or non-EU). Consumers can seek compensation in complex cases, including safety/security regulation breaches (the draft AI Liability Directive (AILD [19]) further specifies such motions in the context of AI software [^2] AILD draft revoked: https://www.euractiv.com/section/tech/news/commission-plans-to-withdraw-ai-liability-directive-draw-mixed-reactions/). The removal of arbitrary thresholds allows full compensation for damages. With a very broad definition of product, this PLD addresses new technologies such as cyber vulnerabilities, essential digital services, and software updates. Software components include any item (e.g., a software library), raw material, or service (e.g., when software initiates remote calls to a SaaS instance) integrated into or connected with a product, though standalone services are generally not covered. Notably, free and open-source software (OSS) outside commercial activity is excluded.
Software as (mostly) a service.
@@
In addition to self-attestation, a cybersecurity “safe harbor”, based on the certification of organizations (with a similar approach already codified in the EU NIS2 Directive [23]), has received considerable attention [24]. Proponents argue that a certification safe harbor can (over time) create incentives to improve collective security practices, including supply chain security, by involving certified upstream technology providers. Dempsey further suggests a hybrid approach, combining a product-based “floor” for liability and a process-based “ceiling” for safe harbor [4]. His proposed mechanism for liability is three-fold. First, a rule-based approach would define the minimum standard of care for software, focusing on product features and behaviors to be included or avoided (e.g., any weaknesses tracked in the MITRE Common Weakness Enumeration (CWE) database [^3] https://cwe.mitre.org/). Second, as software can be complex and dynamic, a liability regime and analysis are needed to assess non-trivial design and implementation flaws; essentially, a software audit. This approach shares similarities with European legislation that has already codified voluntary [25] and mandatory [26] software audit and certification. Third, to provide a safe harbor to developers who followed secure coding recommendations, the proposal suggests a liability waiver or limitation, which makes the outcome predictable for vendors. Clearly defining the conditions under which a vendor qualifies for safe harbor provides legal certainty, reducing the need for costly and prolonged expert debates over what qualifies as reasonable (non-negligent) conduct [7]. A novel development in the AI arena is California Senate Bill 813 [27], proposing a type of safe harbor (affirmative defense) for AI developers against certain civil liability claims when systems are certified by approved and independent Multistakeholder Regulatory Organizations (MROs).
@@ 3 Designing A Better Software Ecosystem
Our proposal begins in the upper path, establishing a minimum set of software development requirements (the floor) and safe harbor practices (the ceiling). As these policies are adopted, they increase the liability of commercial software developers for their software products. In addition, a set of process and product audits, combined with other measures, increase transparency by signaling software quality to stakeholders. Together, these forces – liability and transparency – combine to increase the accountability of vendors. This, in turn, increases the cost of releasing low-quality (i.e., vulnerable, untested, unmaintained) software, which in turn reduces the vendor’s incentive to rush to release software.
@@ 3.1 Audit and Software Clearinghouse
The audit service should, at a minimum, include the following components. (1) Completing an audit should, in general, be optional/voluntary, but may be mandatory in some specific and limited scope; e.g., certain types of software, such as those used in critical infrastructure, may be required to go through an audit. (2) Passing an audit waives the vendor of future liability up to some upper limit for fines or penalties it may face if/when vulnerabilities are discovered or if harm is attributed to its software. (3) The audit process itself should go through regular updates and improvements. Thus, the audit also carries a version number that is included in the certification it issues. (4) Software that fails an audit should not be allowed to be released.
@@
One of the key outcomes of the audit system will be a Software Clearinghouse, a repository of software that has been certified by the auditor, including libraries that feed into a myriad of other software products and systems. In principle, software in the clearinghouse has been certified, and its subsequent use (e.g., as dependencies, clearly articulated in the software’s Bill of Materials (SBOM) [28]) carries a limited liability waiver as described in the previous section. We believe this will encourage customers to maximally use certified software and vendors to build on already certified dependencies. Such an approach has the potential to (partially) prevent the challenges and well-documented struggles of customers trying to identify vulnerable dependencies within their own systems upon the disclosure of newly identified vulnerabilities (especially affecting popular libraries/packages), just as it happened in the Log4Shell case [1].
@@
Software product and process audits have their own set of challenges; some even question whether cybersecurity-related audits could be effective [29]. In reality, software quality assurance in general and audit, in particular, is its own field of study [30] with its own standards and recommendations, e.g., IEEE 1028-2008 [31]. In relation to product-based security testing, defining a set of minimum requirements (i.e., a “floor”) can enhance both the effectiveness of the audit and the interpretation of the resulting certification (see Section 3.2 for details). A “simple” audit similar to enforcing “building codes” in the construction industry could be an adequate starting point [32].
@@
As for the technical security testing methodology, the widespread DevOps philosophy and the resulting Continuous Integration/Continuous Development (CI/CD) approach, designed to streamline and speed up software production and feature delivery, pose unique challenges. One such challenge is the need for rapid and continuous security testing, manifesting in the DevSecOps trend. Such speed-up requires automation, which is far from trivial for more powerful testing methods, e.g., dynamic testing [33]. In the context of the proposed Software Clearinghouse, one critical enabling technology will be quick and automated (regression) testing, to keep up with the rapid releases of new versions of already certified software packages without re-doing a full audit [34].
@@ 3.2 Software Liability
This approach can directly lead to a meaningful interpretation of liability, as it delineates the responsibility: the vendor is (largely) not responsible for the residual risk its product carries as long as it adheres to a specified standard of process; it is, however, wholly responsible for meeting a set of minimum quality requirements even if it has adhered to the process standards. From the user’s point of view, this delineation helps them understand where their own responsibilities and risks lie: if a software has passed both the product audit and the process audit, then the user is aware that the vendor can no longer be faulted for harm that may be caused by the software. If, on the other hand, the software passed the product audit but failed the process audit, then the vendor may be responsible for compensating the user for a large share of the harm attributable to the software. In either case, the user is better informed and can use that knowledge to seek the right type of risk protection (see Section 3.3 on insurance). In the same vein, this approach can help resolve legal disputes by establishing who is responsible when harm occurs.
@@
As to how to define the minimum quality requirements, i.e., the floor, one may turn to “unforgivable vulnerabilities” [35]. Such weaknesses are trivial to find and occur time and time again, demonstrating a “systematic disregard” for secure development practices. In a recent proposal, the UK National Cyber Security Centre (NCSC) proposed a method to classify vulnerabilities into “forgivable” and “unforgivable”  [36]. The method focuses on finding the root cause of vulnerabilities and their respective top-level mitigation based on the MITRE CWE catalog. The mitigation is then assigned an “ease of implementation” score (taking cost, familiarity, and technical feasibility into consideration); vulnerabilities with an “easy” mitigation are then declared unforgivable. Simply put, such weaknesses should not appear in reasonably secure software.
@@
In turn, the process-based ceiling may adopt ideas from the secure software development lifecycle, e.g., Microsoft’s SDL [37]. SDL promotes a DevSecOps approach, integrating security into DevOps (DevSecOps). It applies to all software development models and supports diverse software types (firmware, AI, IoT, web services, etc.) across multiple platforms (cloud, on-premises, mobile, SaaS, etc.). The SDL emphasizes 10 key security practices to enhance software security throughout the whole lifecycle. Adding to this, the 2024 White House report on secure and measurable software [38] further promotes the use of advanced secure technologies such as memory-safe programming languages (e.g., Rust, Go, etc.), memory-safe hardware, and formal methods to prove the correctness of the developed code. Furthermore, the report advocates for the use of cybersecurity quality metrics. Fascinatingly, it also acknowledges the need for shifting market forces, i.e., realigning incentives, to improve these metrics, which is exactly our objective in this paper.
@@
Interestingly, the floor-ceiling approach can also help settle the two key pending legal issues concerning software liability mentioned earlier: (1) whether software will remain a “service” in the eyes of the law or will start being viewed as a “product” at some point; and (2) whether legal protections against lawsuits should be based on satisfactory process, product quality, or both. As far as we know, there have been no direct liability lawsuits until very recently (see Delta Airlines vs. CrowdStrike in our case study in Section 4). This is directly tied to the fact that there is a lack of more precise definitions of liability, which is much needed. As a reference, the EU GDPR was able to define both duty and standard of care, leading to substantial fines for mishandling personal data and prompting a reassessment of cybersecurity investments (as risk-matching cybersecurity countermeasures have also been prescribed for safeguarding personal data). This gives one hope that similar definitions may emerge for the software industry.